# Multisymplectic observable reduction using constraint triples

### v1

@@ Introduction
Reduction can be described, in very broad strokes, as a procedure procedure that associates to a given object (such as a manifold or an algebra) a “smaller” object of the same kind. The main mechanisms by which this can happen are by looking at subspaces or by dividing out symmetries, that is, taking quotients. In the context of mechanical systems admitting some version of the Noether theorem, i.e. that have a link between symmetries and conserved quantities, both mechanisms are usually applied at the same time, in the form of a suboject-quotient reduction scheme.
@@
In this article, we provide a conceptual explanation for the results of [BMR24] by interpreting them using the constraint triples originally introduced by Dippel, Esposito and Waldmann [DEW19] as an algebraic framework for coisotropic reduction. In doing so, our constructions become very general: We obtain a procedure for creating (and reducing) an $L_{\infty}$-algebra (of observables) given any Gerstenhaber algebra equipped with a BV-module and a cocycle.
@@
Under these hypotheses, the MWM theorem asserts that the quotient $M_{\textrm{red}}=\hat{\mu}^{-1}(0)\sslash G$ is a smooth manifold that inherits a uniquely determined symplectic form $\omega_{\textrm{red}}$ such that $\pi^{*}\omega_{\textrm{red}}=i^{*}\omega$, where $\pi:\hat{\mu}^{-1}(0)\to M_{\textrm{red}}$ is the canonical projection and $i:\hat{\mu}^{-1}(0)\hookrightarrow M$ is the inclusion map. Such a result has several far-reaching consequences. It provides, for instance, the symplectic structure on coadjoint orbits, realised as reductions of the cotangent bundle $T^{*}G$ under the conjugation action of $G$ on itself. More generally, it underlies cotangent bundle reduction: given a coWheneverion space $Q$ with a free and proper $G$-action, the canonical Hamiltonian $G$-action on the phase space $M=T^{*}Q$ admits a momentum map, and the reduced space at zero momentum is symplectomorphic to $T^{*}(Q\sslash G)$. In other words, the theorem provides a geometric formulation of Noether’s principle: Whenever a Hamiltonian system with symmetry group $G$ admits a momentum map, the conserved quantities associated with the symmetries are encoded in its image. Moreover, the dynamics of the system descends to the reduced phase space $M\sslash G$, obtained via symplectic reduction.
@@
The key observation is that a general reduction scheme can be formalized as a subobject–quotient process: starting from a total space $M_{\scriptscriptstyle T}$, one first restricts to a subspace $M_{\scriptscriptstyle N}\subseteq M_{\scriptscriptstyle T}$, and then imposes an equivalence relation $\sim_{\scriptscriptstyle 0}$ on $M_{\scriptscriptstyle N}$, obtaining the reduced space $M_{\scriptscriptstyle 0}:=M_{\scriptscriptstyle N}/\sim_{\scriptscriptstyle 0}$. For instance, in classical MWM reduction, the total space is a symplectic manifold $M_{\scriptscriptstyle T}=M$, the subspace is the zero level set $M_{\scriptscriptstyle N}=\mu^{-1}(0)$ of the moment map, and the equivalence relation is induced by the Lie group action, i.e. $\sim_{\scriptscriptstyle 0}=\sim_{G}$.
@@
$\mathop{}\!{\mathrm{d}}\colon V^{i}\longrightarrow V^{i+1},\quad d^{2}=0.$
@@ Acknowledgements
The authors thank Casey Blacker, Peter Crooks, Marco Valerio D’Agostino, Marvin Dippel, Chiara Esposito, Domenico Fiorenza, David Kern, Stefan Waldmann, and Luca Vitagliano for their many fruitful discussions, which contributed significantly to the development of this manuscript. LR was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Projektnummer 539126009 and acknowledges the financial support of the LabEx MiLyon. He thanks Sapienza University of Rome for its hospitality during his research stay.
@@ 1.1 Symmetries and reduction
satisfying $\pi_{\mathfrak{X}(M)}(\mu(\xi))=\underline{\xi}$ for all $\xi\in\mathfrak{g}$.
We will, by abuse of notation, identify $\mu$ with its composition $\pi_{\Omega(M)}\circ\mu$, since the infinitesimal action always determines the vector field component.
@@
  • A covariant momentum map $\mu$, which is $G$-invariant when regarded as a map $\hat{\mu}:M\to\Lambda^{k}T^{*}M\boxtimes\mathfrak{g}^{*}$.
@@ 3.2 Constraint BV-modules
Let $\mathfrak{L}$ be a constraint Lie–Rinehart algebra over a constraint algebra $A$. Then the constraint vector space $\Lambda_{A}\mathfrak{L}$ is constraint Gerstenhaber algebra and $\mathrm{CE}(\mathfrak{L})=((\Lambda_{A}\mathfrak{L})^{*})^{\mathrm{rev}}$ is a constraint BV-module over it.
@@ 4.1 Constraint $L_{\infty}$-algebras and their reduction
By its very definition, $\mathsf{Ham}(V,\omega)_{\scriptscriptstyle T}$ is an $L_{\infty}$-algebra since is given specializing Subsection 2.2 to the BV-module $V_{\scriptscriptstyle T}$ over $G_{\scriptscriptstyle T}$. Thus, we only have to check that the brackets are constraint morphisms. For this, we consider $l_{1}$ and the remaining cases separately:
  • Since $B$ is a constraint BV-module, $l_{1}:\mathsf{Ham}^{i}(V,\omega)\to\mathsf{Ham}^{i+1}(V,\omega)$ is a constraint morphism for $i<-1$. For $i=1$, the same holds true because the additional component 00 in the expression $l_{1}(\alpha)=(d\alpha,0)$ lies in $V_{\scriptscriptstyle 0}$.
@@ 4.2 Constraint $L_{\infty}$-algebra induced by symmetries
We now interpret the constraint Lie–Rinehart algebra $(A^{\prime},\mathfrak{Y})$ as a subquotient of $(A,\mathfrak{X}(A))$: that is, it can be obtained by first restricting to a subalgebra and then passing to a quotient. Concretely, this corresponds to the diagram
$\left(\begin{array}[]{l}\mathfrak{X}(A)_{\scriptscriptstyle T}\\ \mathfrak{X}(A)_{\scriptscriptstyle N}\\ \mathfrak{X}(A)_{\scriptscriptstyle 0}\end{array}\right)\longleftarrow\left(\begin{array}[]{l}\mathfrak{X}(A)_{\scriptscriptstyle T}\\ \mathfrak{Y}_{\scriptscriptstyle N}\\ \mathfrak{X}(A)_{\scriptscriptstyle 0}\end{array}\right)\longrightarrow\left(\begin{array}[]{l}\mathfrak{Y}_{\scriptscriptstyle T}\\ \mathfrak{Y}_{\scriptscriptstyle N}\\ \mathfrak{Y}_{\scriptscriptstyle 0}\end{array}\right).$
@@
    $\mathcal{L}_{X}d\alpha$ $=\mathop{}\!{\mathrm{d}}\mathcal{L}_{X}\alpha\in\mathop{}\!{\mathrm{d}}B_{\scriptscriptstyle 0}\subseteq B_{\scriptscriptstyle 0}$
@@ 4.4.2 Multisymplectic reduction with respect to a covariant momentum map
Let $(M,\omega)$ be a multisymplectic manifold with the action of a connected Lie group $G$, and $\mu$ a covariant moment map for this action (see Subsection 1.1), such that its vanishing locus $S$ is an embedded submanifold and the action of $G$ on $S$ is free and proper. Then we can apply [Bla21, Theorem 1], to obtain a pre-mulisymplectic manifold $(M_{\textrm{red}},\omega_{\textrm{red}})$, where $M_{\textrm{red}}={S}\sslash{G}$ and $\omega_{\textrm{red}}$ is compatible with $\omega$ when pulled back to $S$. We note that, since $G$ is connected, its fundamental vector fields form a regular foliation when restricted to $S$; in particular, we are in the setting of the previous subsection.

### v2

@@ Introduction
Reduction can be described, in very broad strokes, as a procedure that associates to a given object (such as a manifold or an algebra) a “smaller” object of the same kind. The main mechanisms by which this can happen are by looking at subspaces or by dividing out symmetries, that is, taking quotients. In the context of mechanical systems admitting some version of the Noether theorem, i.e. that have a link between symmetries and conserved quantities, both mechanisms are usually applied at the same time, in the form of a subobject-quotient reduction scheme.
@@
In this article, we provide a conceptual explanation for the results of [BMR24] by interpreting them using the constraint triples originally introduced by Dippell, Esposito and Waldmann [DEW19] as an algebraic framework for coisotropic reduction. In doing so, our constructions become very general: We obtain a procedure for creating (and reducing) an $L_{\infty}$-algebra (of observables) given any Gerstenhaber algebra equipped with a BV-module and a cocycle.
@@
Under these hypotheses, the MWM theorem asserts that the quotient $M_{\textrm{red}}=\hat{\mu}^{-1}(0)/G$ is a smooth manifold that inherits a uniquely determined symplectic form $\omega_{\textrm{red}}$ such that $\pi^{*}\omega_{\textrm{red}}=i^{*}\omega$, where $\pi:\hat{\mu}^{-1}(0)\to M_{\textrm{red}}$ is the canonical projection and $i:\hat{\mu}^{-1}(0)\hookrightarrow M$ is the inclusion map. Such a result has several far-reaching consequences. It provides, for instance, the symplectic structure on coadjoint orbits, realised as reductions of the cotangent bundle $T^{*}G$ under the conjugation action of $G$ on itself. More generally, it underlies cotangent bundle reduction: given a configuration space $Q$ with a free and proper $G$-action, the canonical Hamiltonian $G$-action on the phase space $M=T^{*}Q$ admits a momentum map, and the reduced space at zero momentum is symplectomorphic to $T^{*}(Q/G)$. In other words, the theorem provides a geometric formulation of Noether’s principle: Whenever a Hamiltonian system with symmetry group $G$ admits a momentum map, the conserved quantities associated with the symmetries are encoded in its image. Moreover, the dynamics of the system descends to the reduced phase space $M/G$, obtained via symplectic reduction.
@@
The key observation is that a general reduction scheme can be formalized as a subobject–quotient process: starting from a total space $M_{\scriptscriptstyle T}$, one first restricts to a subspace $M_{\scriptscriptstyle N}\subseteq M_{\scriptscriptstyle T}$, and then imposes an equivalence relation $\sim_{\scriptscriptstyle 0}$ on $M_{\scriptscriptstyle N}$, obtaining the reduced space $M_{\textrm{red}}:=M_{\scriptscriptstyle N}/\sim_{\scriptscriptstyle 0}$. For instance, in classical MWM reduction, the total space is a symplectic manifold $M_{\scriptscriptstyle T}=M$, the subspace is the zero level set $M_{\scriptscriptstyle N}=\mu^{-1}(0)$ of the moment map, and the equivalence relation is induced by the Lie group action, i.e. $\sim_{\scriptscriptstyle 0}=\sim_{G}$.
@@
$\mathop{}\!{\mathrm{d}}\colon V^{i}\longrightarrow V^{i+1},\quad\mathop{}\!{\mathrm{d}}^{2}=0.$
@@ Acknowledgements
The authors thank Casey Blacker, Peter Crooks, Marco Valerio D’Agostino, Marvin Dippell, Chiara Esposito, Domenico Fiorenza, David Kern, Stefan Waldmann, and Luca Vitagliano for their many fruitful discussions, which contributed significantly to the development of this manuscript. The authors also thank the editor and referee for their careful reading of the manuscript. LR was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Projektnummer 539126009 and acknowledges the financial support of the LabEx MiLyon. He thanks Sapienza University of Rome for its hospitality during his research stay.
@@ 1.1 Symmetries and reduction
satisfying $\pi_{\mathfrak{X}(M)}(\mu(\xi))=\underline{\xi}$ for all $\xi\in\mathfrak{g}$, where $\pi_{\mathfrak{X}(M)}$ is the projection $\mathsf{Ham}^{0}(M,\omega)\subseteq\Omega^{k-1}(M)\times\mathfrak{X}(M)\to\mathfrak{X}(M)$.
We also write $\pi_{\Omega(M)}$ for the projection $\mathsf{Ham}^{0}(M,\omega)\subseteq\Omega^{k-1}(M)\times\mathfrak{X}(M)\to\Omega^{k-1}(M)$. Moreover, we will, by abuse of notation, identify $\mu$ with its composition $\pi_{\Omega(M)}\circ\mu$, since the infinitesimal action always determines the vector field component.
@@
  • A covariant momentum map $\mu$, which is $G$-invariant when regarded as a map $\hat{\mu}:M\to\Lambda^{k}T^{*}M\boxtimes\mathfrak{g}^{*}$, where $\boxtimes$ denotes the external tensor product.
@@ 3.2 Constraint BV-modules
Let $\mathfrak{L}$ be a constraint Lie–Rinehart algebra over a constraint algebra $A$. Then the constraint vector space $\Lambda_{A}\mathfrak{L}$ is a constraint Gerstenhaber algebra and $\mathrm{CE}(\mathfrak{L})=((\Lambda_{A}\mathfrak{L})^{*})^{\mathrm{rev}}$ is a constraint BV-module over it.
@@ 4.1 Constraint $L_{\infty}$-algebras and their reduction
By its very definition, $\mathsf{Ham}(V,\omega)_{\scriptscriptstyle T}$ is an $L_{\infty}$-algebra since it is given specializing Subsection 2.2 to the BV-module $V_{\scriptscriptstyle T}$ over $G_{\scriptscriptstyle T}$. Thus, we only have to check that the brackets are constraint morphisms. For this, we consider $l_{1}$ and the remaining cases separately:
  • Since $V$ is a constraint BV-module, $l_{1}:\mathsf{Ham}^{i}(V,\omega)\to\mathsf{Ham}^{i+1}(V,\omega)$ is a constraint morphism for $i<-1$. For $i=1$, the same holds true because the additional component 00 in the expression $l_{1}(\alpha)=(d\alpha,0)$ lies in $V_{\scriptscriptstyle 0}$.
@@ 4.2 Constraint $L_{\infty}$-algebra induced by symmetries
We now interpret the constraint Lie–Rinehart algebra $(A^{\prime},\mathfrak{Y})$ as a subquotient of $(A,\mathfrak{X}(A))$: that is, it can be obtained by first restricting to the subalgebra and then passing to a quotient. Concretely, this corresponds to the following diagram which should be read from right to left.
$\left(\begin{array}[]{l}\mathfrak{X}(A)_{\scriptscriptstyle T}\\ \mathfrak{X}(A)_{\scriptscriptstyle N}\\ \mathfrak{X}(A)_{\scriptscriptstyle 0}\end{array}\right)\overset{\mathrm{quotient}}{\longleftarrow}\left(\begin{array}[]{l}\mathfrak{X}(A)_{\scriptscriptstyle T}\\ \mathfrak{Y}_{\scriptscriptstyle N}\\ \mathfrak{X}(A)_{\scriptscriptstyle 0}\end{array}\right)\overset{\mathrm{subspace}}{\longrightarrow}\left(\begin{array}[]{l}\mathfrak{Y}_{\scriptscriptstyle T}\\ \mathfrak{Y}_{\scriptscriptstyle N}\\ \mathfrak{Y}_{\scriptscriptstyle 0}\end{array}\right).$
@@
    $\mathcal{L}_{X}\mathop{}\!{\mathrm{d}}\alpha$ $=\mathop{}\!{\mathrm{d}}\mathcal{L}_{X}\alpha\in\mathop{}\!{\mathrm{d}}B_{\scriptscriptstyle 0}\subseteq B_{\scriptscriptstyle 0}$
@@ 4.4.2 Multisymplectic reduction with respect to a covariant momentum map
Let $(M,\omega)$ be a multisymplectic manifold with the action of a connected Lie group $G$, and $\mu$ a covariant moment map for this action (see Subsection 1.1), such that its vanishing locus $S$ is an embedded submanifold and the action of $G$ on $S$ is free and proper. Then we can apply [Bla21, Theorem 1], to obtain a pre-mulisymplectic manifold $(M_{\textrm{red}},\omega_{\textrm{red}})$, where $M_{\textrm{red}}={S}/{G}$ and $\omega_{\textrm{red}}$ is compatible with $\omega$ when pulled back to $S$. We note that, since $G$ is connected, its fundamental vector fields form a regular foliation when restricted to $S$; in particular, we are in the setting of the previous subsection.